# Partial Wavefunction Collapse Under Repeated Weak Measurement of a non-Conserved Observable

### v1

@@ IV Conclusion and Outlook
Our key result is that repeated weak measurements of a non-conserved observable can extract information about other observables and lead to a partial collapse of the system’s density matrix. On one hand, this is intriguing because it sheds more light on the measurement problem and the nature of wavefunction collapse, reinforcing the idea that a quantum system tends towards collapse whenever information is extracted from it, even if such information is gained indirectly. On the other hand, this result also presents a new approach to measurement that may have some utility in experiment. For example, if an observable of interest is difficult to measure directly it may be possible to find a different observable which is easier to measure yet still indirectly extracts the desired information. This is highlighted in the Heisenberg chain example, where measurement of a local observable was able to reveal information about a total system observable. In particular, such a protocol may find a useful application in Rydberg array experiments where the use of ancilla qubits coupled to one or a few sites is very natural but performing certain direct measurements may not be.

### v2

@@ IV Conclusion and Outlook
Our key result is that repeated weak measurements of a non-conserved observable can extract information about other observables and lead to a partial collapse of the system’s density matrix, giving rise to a new potential class of measurement protocols we term Auxiliary Observable QND. On one hand, this is intriguing because it sheds more light on the measurement problem and the nature of wavefunction collapse, reinforcing the idea that a quantum system tends towards collapse whenever information is extracted from it, even if such information is gained indirectly. On the other hand, this result also presents a new approach to measurement that may have some utility in experiment. For example, if an observable of interest is difficult to measure directly it may be possible to find a different observable which is easier to measure yet still indirectly extracts the desired information. This is highlighted in the Heisenberg chain example, where measurement of a local observable was able to reveal information about a total system observable. In particular, such a protocol may find a useful application in Rydberg array experiments where coupling ancilla qubits to one or a few sites is a very natural operation, so this type of ancilla-based indirect measurement may be easier to implement than a direct measurement.
Acknowledgements.
This work was supported by National Science Foundation Division of Mathematical Sciences Grant No. 2138905.